### v1

# When is the Carathéodory metric a Kähler metric?

Abstract
We prove that if the Carathéodory metric on a strictly pseudoconvex domain with a smooth boundary is a Kähler metric, then the domain is biholomorphic to a ball. Additionally, we highlight the relationships between the Lu constant and the holomorphic Bergman curvature.
footnotetext: Keywords. Carathéodory metric, Bergman metric, Kähler metric, Kähler-Einstein metric, biholomorphic map, holomorphic sectional curvature, Lu constant, strictly pseudoconvex domain Mathematics Subject Classification. Primary 32F45; Secondary 32T15, 32H02
1 Introduction
A bounded pseudoconvex domain in complex Euclidean space possesses several biholomorphically invariant metrics: the Bergman metric, the differential Carathéodory metric, the unique (up to scaling) Kähler-Einstein metric, and the differential Kobayashi-Royden metric. These metrics all coincide, up to a multiplicative constant, on the unit ball $\mathbb{B}^{n}$. However, this coincidence does not generally occur for arbitrary domains, leading to the following natural question:
Question 1.1.
What are the domains where the Bergman, Carathéodory, Kähler-Einstein, and Kobayashi metrics are not pairwise distinct (up to scaling)?
Question 1.1 relates to a general problem raised by Yau [45, pp. 679] (in a slightly different form) on the characterization of manifolds whose Bergman metrics are Kähler-Einstein. Cheng and Yau [9] showed the existence of a complete Kähler-Einstein metric on a smoothly bounded strictly pseudoconvex domain in $\mathbb{C}^{n}$. Mok and Yau [33] later demonstrated that there is a complete Kähler-Einstein metric for any bounded pseudoconvex domain without boundary regularity conditions. For smoothly bounded strictly pseudoconvex domains, Cheng conjectured that the Bergman metric is Kähler-Einstein if and only if the domain is biholomorphic to a ball. Following the earlier works of Fu and the third author [19] and Nemirovski and Shafikov [36], this conjecture was confirmed by Huang and Xiao in [25]. Additionally, see the recent work [38] by Savale and Xiao for finite type domains in $\mathbb{C}^{2}$.
A famous theorem of Lu [31] says that a bounded domain in $\mathbb{C}^{n}$ with a complete Bergman metric of constant holomorphic sectional curvature must be biholomorphic to $\mathbb{B}^{n}$. In [16, 17], the first and third authors extended Lu’s theorem to the Bergman-incomplete situation and provided a multidimensional generalization of Carathéodory’s theorem on the continuous extension of biholomorphisms up to the closures. See also the recent work [24] by Huang and Li.
For a bounded domain $\Omega$ in $\mathbb{C}^{n}$, denote by $\mathbb{D}(\Omega)$ the family of holomorphic functions from $\Omega$ into $\mathbb{D}$, and denote by $\Omega(\mathbb{D})$ the family of holomorphic mappings from $\mathbb{D}$ into $\Omega$, where $\mathbb{D}$ denotes the unit disk in $\mathbb{C}$. Recall (see [43]) that:
  1. $1$
    The differential Carathéodory metric is given by
    $\begin{split}&C_{\Omega}:\Omega\times\mathbb{C}^{n}\to\mathbb{R}^{+}\cup\{0\},\\ &C_{\Omega}(z,v)=\sup\{|df(v)|:f\in\mathbb{D}(\Omega),f(z)=0\},\end{split}$
    where $df(v)$ is measured with respect to the Poincaré metric on $\mathbb{D}$;
  2. $2$
    The differential Kobayashi metric is given by
    $\begin{split}&K_{\Omega}:\Omega\times\mathbb{C}^{n}\to\mathbb{R}^{+}\cup\{0\},\\ &K_{\Omega}(z,v)=\inf\{|t|:t\text{ is a tangent vector to }\mathbb{D}\text{ at}0,\\ &\exists f\in\Omega(\mathbb{D})\text{ with }f(0)=z\text{ and }df(t)=v\},\end{split}$
    where $t$ is measured with respect to the Poincaré metric on $\mathbb{D}$;
In the seventies, the third author observed in [42] that on a complex manifold $M$, the holomorphic sectional curvature of the differential Kobayashi metric $K_{M}$ is bounded below by $-2$, and that of the differential Carathéodory metric $C_{M}$ is bounded above by $-2$. With this observation, one can immediately conclude the following characterization of the unit ball.
Theorem 1.2 ([42]).
Let $M$ be a simply-connected complete hyperbolic complex manifold with $C_{M}=K_{M}$, assumed to be a Kähler metric. Then, $M$ is biholomorphic to a ball.
A Kähler manifold is a complex manifold with a Hermitian metric whose associated 2-form is closed. Stanton proved the following result, which is more general than Theorem 1.2.
Theorem 1.3 ([40]).
Let $M$ be a complete hyperbolic complex manifold. Assume that there is a point on $M$ at which $C_{M}$ and $K_{M}$ are equal, and one of these metrics is Hermitian and of class $C^{\infty}$. Then $M$ is biholomorphic to a ball.
In [10, 11], Cheung and the third author proved that if a bounded convex domain in $\mathbb{C}^{n}$ has a Hermitian metric of constant negative holomorphic sectional curvature, such that all metric components blow up to infinity at the boundary, then the domain is biholomorphic to $\mathbb{B}^{n}$.
In this note, we investigate the bounded domains in $\mathbb{C}^{n}$ where the Carathéodory metric coincides up to scaling with the Bergman or Kähler-Einstein metric, or more generally, is a Kähler metric. Our work is inspired by a recent paper by Gaussier and Zimmer [20], who used the Kobayashi metric to characterize domains whose universal coverings are biholomorphic to a ball (see the theorem below); they also obtained results concerning locally symmetric metrics.
Theorem 1.4 ([20]).
Suppose that $\Omega\subset\mathbb{C}^{n}$, $n\geq 2$, is a bounded strictly pseudoconvex domain with a $C^{2}$ boundary. Then, the following are equivalent:
  1. $(a)$
    The Kobayashi metric on $\Omega$ is a Kähler metric;
  2. $(b)$
    The Kobayashi metric on $\Omega$ is a Kähler metric with constant holomorphic sectional curvature;
  3. $(c)$
    The universal cover of $\Omega$ is biholomorphic to a ball;
  4. $(d)$
    The Kobayashi metric on $\Omega$ is a Kähler-Einstein metric.
Moreover, the Kobayashi metric on $\Omega$ is a scalar multiple of the Bergman metric if and only if $\Omega$ is biholomorphic to a ball.
For strictly pseudoconvex domains, we obtain the following main result, which is a Carathéodory-analogue of both Lu’s uniformization theorem and the resolution of the Cheng conjecture.
Theorem 1.5.
Suppose that $\Omega\subset\mathbb{C}^{n}$, $n\geq 2$, is a bounded strictly pseudoconvex domain with a $C^{2}$ boundary. Then, the following are equivalent:
  1. $(1)$
    The Carathéodory metric on $\Omega$ is a Kähler metric;
  2. $(2)$
    The Carathéodory metric on $\Omega$ is a Kähler metric with constant holomorphic sectional curvature;
  3. $(3)$
    $\Omega$ is biholomorphic to the unit ball;
  4. $(4)$
    The Carathéodory metric on $\Omega$ is a scalar multiple of the Bergman metric;
  5. $(5)$
    The Carathéodory metric on $\Omega$ is a Kähler-Einstein metric.
For $n=1$, Suita [41] (cf. [4]) showed that the Carathéodory metric is real analytic for any bounded domain in $\mathbb{C}$. Thus, it can be easily proven that if the Carathéodory metric is complete with holomorphic sectional curvature identically equal to $-2$, then the domain is biholomorphic to a disc.
Some related examples are discussed below.
Examples.
  1. $(i)$
    By the remarkable works of Lempert [28, 27, 29], the Carathéodory metric and the Kobayashi metric coincide on convex domains. Theorem 1.2/1.3 implies that the Carathéodory metric on a bounded convex domain is a Hermitian metric if and only if the domain is biholomorphic to the unit ball. In particular, let $\mathcal{P}$ be a small generic perturbation of $\mathbb{B}^{n}$ such that $\mathcal{P}$ is convex in $\mathbb{C}^{n}$ but not biholomorphic to $\mathbb{B}^{n}$ when $n\geq 2$ (see the papers of Burns, Shnider and Wells [7], and Greene and Krantz[22]). Then, the Carathéodory metric on $\mathcal{P}$ is not a Hermitian metric when $n\geq 2$.
    In [14], the first author and Treuer constructed a bounded complete Reinhardt domain
    :=$\mathcal{R}:=\{z\in\mathbb{C}^{2}:|z_{1}|^{4}+|z_{1}|^{2}+|z_{2}|^{2}<1\}$
    which is strongly convex with an algebraic boundary and not biholomorphic to $\mathbb{B}^{2}$. By the above discussion, the Carathéodory metric on $\mathcal{R}$ is not a Hermitian metric.
    Also, note that a complex analytic ellipsoid
    ${E}:=\{(z,w)\in{\mathbb{C}^{1}}\times{\mathbb{C}^{1}}:|z|^{2}+|w|^{2p}<1\},$
    is convex and not biholomorphic to $\mathbb{B}^{2}$, where $\mathbb{Z}^{+}\ni p>1$. So, the Carathéodory metric on ${E}$ is not a Hermitian metric. (See some recent related work in this direction by Cho [12]).
  2. $(ii)$
    The Kobayashi metric of the unit bidisc is
    $K_{\mathbb{D}^{2}}((z,w),(X,Y))=\max\{K_{\mathbb{D}}(z,X),K_{\mathbb{D}}(w,Y)\}.$
    Thus, the Carathéodory metric and the Kobayashi metric are not $C^{1}$-smooth metrics in general. For bounded symmetric domains of rank $\geq 2$, the Carathéodory and Kobayashi metrics are not Kähler.
  3. $(iii)$
    Burns and Shnider [6] have constructed examples of non-simply connected strictly pseudoconvex domains with spherical boundaries in $\mathbb{C}^{n}$. By the uniformization theorem (see [34, 35]), such domains are universally covered by the unit ball, so the Kobayashi metric is Kähler-Einstein. For example, let
    $\mathcal{S}:=\{(z,w)\in\mathbb{C}^{2}\mid\sin\log|z|+|w|^{2}<0,e^{-\pi}<|z|<1\}.$
    By our Theorem 1.5 above, the Carathéodory metric $C_{\mathcal{S}}$ is not a Kähler metric on $\mathcal{S}$, and thus $C_{\mathcal{S}}$ is not equal to the Kobayashi metric $K_{\mathcal{S}}$.
The following significant fact, that the Bergman metric always dominates the Carathéodory metric on bounded Euclidean domains, is an immediate consequence of an inequality due to Lu.
Theorem 1.6 ([30], cf. [8]).
Let $\Omega$ be a bounded domain in $\mathbb{C}^{n}$. For each tangent vector $v\in T_{z}(\Omega)=\mathbb{C}^{n}$ at $z\in\Omega$, it holds that
$B_{\Omega}(z,v)\geq C_{\Omega}(z,v),$
where $B_{\Omega}(z,v)$ and $C_{\Omega}(z,v)$ are the lengths of $v$ with respect to the Bergman metric $B^{2}_{\Omega}$ and the differential Carathéodory metric $C_{\Omega}$, respectively.
The above inequality is not sharp. Therefore, we consider the following Lu constant.
Definition. The Lu constant $L(\Omega)$ of a bounded domain $\Omega$ in $\mathbb{C}^{n}$ is defined as
[…]
The definition of $L(\Omega)$ implies that
$L(\Omega)B_{\Omega}\geq C_{\Omega}.$ (1.1)
For the rest of this paper, we shall call (1.1) the sharp Lu’s inequality. The value of the Lu constant $L(\Omega)$ is understood in the following scenarios:
By Theorem 1.6, $L(\Omega)\leq 1$. Also, $L(\Omega)=\frac{1}{\sqrt{n+1}}$ when $\Omega$ is the unit ball, and $L(\Omega)=\sqrt{2n}$ when $\Omega$ is the unit polydisc $\mathbb{D}^{n}$. One further determines the Lu constants of all Cartan classical domains. When $\Omega$ is a bounded strictly pseudoconvex domain in $\mathbb{C}^{n}$, $L(\Omega)\geq\frac{1}{\sqrt{n+1}}$ by the results on the boundary behaviors of $B_{\Omega}$ and $C_{\Omega}$ due to Diederich [18] and Graham [21], respectively. If the holomorphic sectional curvature of the Bergman metric on a bounded strictly pseudoconvex domain in $\mathbb{C}^{n}$ is bounded below by $-\frac{2}{n+1}$, then $L(\Omega)=\frac{1}{\sqrt{n+1}}$ following from the Ahlfors-Schwarz lemma.
Next, along the line of the Wong/Stanton theorem, we present a rigidity result on the comparison between the Bergman metric and the Carathéodory metric.
Theorem 1.7.
Let $X$ be a hyperbolic Riemann surface. Assume that the Bergman metric $B(z)|dz|^{2}$ is complete with holomorphic sectional curvature no less than $-1$, and the Carathéodory metric is written as $c^{2}_{B}(z)|dz|^{2}$, where $z$ is a local coordinate. Then, it holds on $X$ that
${B}\geq 2c^{2}_{B};$
moreover, $X$ is biholomorphic to a disc if and only if there exists some $z_{0}\in X$ such that
$B(z_{0})=2c^{2}_{B}(z_{0}).$
Let $\Omega\subset\mathbb{C}^{n}$ be a domain whose universal cover is biholomorphic to a Euclidean ball. Assume that the Bergman metric $B^{2}_{\Omega}$ is complete on $\Omega$ with holomorphic sectional curvature no less than $-2(L(\Omega))^{2}$. Then, $\Omega$ is biholomorphic to a ball if and only if there exists some $z_{0}\in\Omega$ such that
$L(\Omega)B_{\Omega}(z_{0},v)=C_{\Omega}(z_{0},v),\quad\text{for all\,}\,v\in T_{z_{0}}(\Omega).$
Since a bounded pseudoconvex domain with a spherical boundary in $\mathbb{C}^{n}$, $n\geq 2$, is universally covered by a ball by the uniformization theorem (see [36, Theorem A.2]), we obtain the following corollary relating to the examples of Burns and Shnider [6] using the aforementioned methods:
Corollary 1.8.
Let $\Omega$ be a bounded strictly pseudoconvex domain with spherical boundary in $\mathbb{C}^{n}$, $n\geq 2$. Then $\Omega$ is not biholomorphic to a Euclidean ball if and only if either the holomorphic sectional curvature of the Bergman metric in some direction is strictly less than $-\frac{2}{n+1}$ or the sharp Lu’s inequality (1.1) is strict at all points and for some directions.
Lastly, we extend a previous result [11] by Cheung and the third author from bounded convex domains to strictly pseudoconvex domains.
Theorem 1.9.
Let $\Omega\subset\mathbb{C}^{n}$, $n\geq 2$, be a bounded strictly pseudoconvex domain with a $C^{2}$ boundary and the Lu constant $L(\Omega)$. Suppose the holomorphic sectional curvature of the Bergman metric $B^{2}_{\Omega}$ on $\Omega$ is bounded above by $-2(L(\Omega))^{2}$. Then $\Omega$ is biholomorphic to a ball.
Besides [20], our proofs of Theorems 1.5 and 1.9 depends heavily on the work of Bracci, Fornæss, and Wold [3]. (For some related observations, see an earlier work by Huang [23].) One of the key ideas is that at a point sufficiently close to the boundary of a smoothly bounded strictly pseudoconvex domain, there is a nonempty, open set of directions where the Kobayashi and Carathéodory metrics coincide. We observe that the Carathéodory metric has constant negative holomorphic sectional curvature in these directions, and in fact, has constant negative holomorphic sectional curvature near the boundary, if the metric is Kähler.
The organization of the paper is as follows. In Section 2, we give a proof of our main result, Theorem 1.5. In Section 3, we provide proofs of Theorems 1.7 and 1.9.
2 Proof of the Main Result
[…]
$\delta_{\Omega}(z)=\min\{\|z-x\|:x\in\partial\Omega\}.$
Definition 2.1.
A holomorphic map $\varphi:\mathbb{D}\rightarrow\Omega$ is a complex geodesic with respect to the Kobayashi (or Carathéodory) metric if
$k_{\Omega}(\varphi(z_{1}),\varphi(z_{2}))=k_{\mathbb{D}}(z_{1},z_{2})\quad(\text{or \,}c_{\Omega}(\varphi(z_{1}),\varphi(z_{2}))=c_{\mathbb{D}}(z_{1},z_{2}))$
for all $z_{1},z_{2}\in\mathbb{D}$, where $k_{\Omega}$ (or $c_{\Omega}$) denotes the Kobayashi (or Carathéodory) distance.
Suppose that ${C_{\Omega}(z,v)}={K_{\Omega}(z,v)}$ for $z\in\Omega,\,v\in T_{z}(\Omega)$, where $\Omega$ is a bounded domain in $\mathbb{C}^{n}$. By the definitions of the Carathéodory and Kobayashi metrics, let $\rho:\Omega\to\mathbb{D}$ be a holomorphic function that realizes the Carathéodory metric $C_{\Omega}(z,v)$, and let $\varphi:\mathbb{D}\to\Omega$ be a holomorphic map that realizes the Kobayashi metric $K_{\Omega}(z,v)$. Then, the composition $\rho\circ\varphi$ is an automorphism of $\mathbb{D}$ that fixes the origin. It can be shown that $\varphi:\mathbb{D}\to\Omega$ is a complex geodesic with respect to both the Carathéodory and Kobayashi metrics. Furthermore, the holomorphic sectional curvatures of $C_{\Omega}$ and $K_{\Omega}$ restricted to the image $\varphi(\mathbb{D})$ are both equal to $-2$ everywhere.
When $\Omega$ has a $C^{2}$ boundary and $z\in\overline{\Omega}$ is sufficiently close to $\partial\Omega$, there is a unique point $\pi(z)\in\partial\Omega$ such that $\|z-\pi(z)\|=\delta_{\Omega}(z)$. We let
$P_{z}:\mathbb{C}^{n}\rightarrow T_{\pi(z)}^{\mathbb{C}}\partial\Omega$
denote the (Euclidean) orthogonal projection and let $P_{z}^{\bot}=\text{Id}-P_{z}$. The following result states that complex geodesics exist at base points sufficiently close to the smooth boundary of strictly pseudoconvex domains and in directions sufficiently tangential.
Theorem 2.2 ([3, 20, 23]).
Suppose that $\Omega\subset\mathbb{C}^{n},n\geq 2,$ is a bounded strictly pseudoconvex domain with $C^{2}$ boundary. Then there exists $\epsilon>0$ such that
${C_{\Omega}(z,v)}={K_{\Omega}(z,v)},\quad z\in\Omega,\,v\in T_{z}(\Omega)$ (2.1)
if $\delta_{\Omega}(z)<\epsilon$ and if
$\left\lVert P_{z}^{\bot}(v)\right\rVert<\epsilon\left\lVert P_{z}(v)\right\rVert.$
Note that (2.1) holds due to the existence of a complex geodesic $\varphi:\mathbb{D}\rightarrow\Omega$ with $\varphi(\mathbb{D})\subset V$, $\varphi(0)=z$, and $\varphi^{\prime}(0)\lambda=v$ for some $\lambda>0$, and a holomorphic map $\rho:\Omega\rightarrow\mathbb{D}$ with $\rho\circ\varphi=\text{id}_{\mathbb{D}}$. When $\Omega$ has a $C^{3}$ boundary, Theorem 2.2 was obtained by Bracci, Fornæss, and Wold [3]; in this case, the existence of $\varphi$ was given by Huang in [23, Theorem 1], and the existence of $\rho$ was given in [3, Theorem 1.1]. See also [5, 26] for related results. Gaussier and Zimmer [20] provided a proof of Theorem 2.2 for the case when $\Omega$ has a $C^{2}$ boundary. They also proved the following result (see Corollary 1.2 and its proof in [20]), which will play a crucial role in our arguments.
Theorem 2.3 ([20]).
Suppose that $M$ is a Stein manifold with $\dim_{\mathbb{C}}M\geq 2$ and $g$ is a complete Kähler metric on $M$. If there exists a compact set $\Sigma\subset M$ such that $g$ has constant negative holomorphic sectional curvature on $M\setminus\Sigma$, then $g$ has constant negative holomorphic sectional curvature on $M$ and the universal cover of $M$ is biholomorphic to a ball in $\mathbb{C}^{\dim_{\mathbb{C}}M}$.
Seshadri and Verma established a version of Theorem 2.3 for simply connected non-positively curved Kähler manifolds in [39].
Proof of Theorem 1.5.
Since $(2)$ implies $(1)$ by definition, and it is well known that $(3)$ implies $(2)$, we only need to prove the non-trivial part, namely that $(1)$ implies $(3)$. By Theorem 2.2, there exists some $\epsilon>0$ such that: if $z\in\Omega$ and $\delta_{\Omega}(z)<\epsilon$, then there is a nonempty, open set $E_{\Omega}(z)$ of directions where the Carathéodory metric and the Kobayashi metric coincide. Using the argument in [42], we know that the holomorphic sectional curvature $H(z;v)$ of the Carathéodory metric is equal to $-2$ for any non-zero $v\in E_{\Omega}(z)$. Briefly speaking, given that the Carathéodory metric and Kobayashi metric coincide at $z$ for $v$, there exists a complex geodesic with respect to the Carathéodory metric on which the holomorphic sectional curvature is equal to $-2$ at every point. As the curvature $H(z;v)$ of $\Omega$ is bounded above by $-2$, the curvature-decreasing property implies that it must be exactly $-2$ at $z$. Moreover, since the holomorphic sectional curvature function
$v\in T_{z}(\Omega)\mapsto H(z;v)\in\mathbb{R}$
is algebraic (with $z$ fixed), it follows that the holomorphic sectional curvature $H(z;v)$ equals $-2$ for all non-zero vectors $v\in T_{z}(\Omega)$. This is true for all $z\in\Omega$ such that $\delta_{\Omega}(z)<\epsilon$ for a sufficiently small $\epsilon$. Define the compact set
$\Sigma:=\{z\in\Omega:\delta_{\Omega}(z)\geq\epsilon\}$ (2.2)
as in Theorem 2.3. Then, the holomorphic sectional curvature of the Carathéodory metric $C_{\Omega}$ is equal to $-2$ on $\Omega\setminus\Sigma$.
By the work of Graham [21, Proposition 5], the Carathéodory metric $C_{\Omega}$ is complete on $\Omega$. If $C_{\Omega}$ is a Kähler metric on $\Omega$, then by Theorem 2.3, the Carathéodory metric $C_{\Omega}$ has constant negative holomorphic sectional curvature on $\Omega$, and the universal cover of $\Omega$ is biholomorphic to the unit ball. By Theorem 1.4, the Kobayashi metric $K_{\Omega}$ is also a complete Kähler metric of constant negative holomorphic sectional curvature on $\Omega$. Since holomorphic covering maps of bounded domains are local isometries for the Kobayashi metric (which is Kähler-Einstein metric on the Euclidean ball), and both $K_{\Omega}$ and $C_{\Omega}$ are complete Kähler metrics of the same constant negative holomorphic sectional curvature on $\Omega$, we see from the Ahlfors-Schwarz lemma (see [2, 37, 44]) that $K_{\Omega}=C_{\Omega}$ on $\Omega$. By Theorem 1.3, one further concludes that $\Omega$ is biholomorphic to a ball.
Regarding Conditions $(4)$ and $(5)$, notice that Condition $(3)$ implies both of them. And either Condition $(4)$ or Condition $(5)$ would imply Condition $(1)$ by definition. Thus, the proof is complete.
Remark. On April 19th, 2024, the first author gave an online talk at Syracuse University’s Analysis Seminar hosted by Yuan Yuan. During the talk, Yuan mentioned that he had obtained some related result yet to be written up. The results in this paper were obtained around 2022/2023.
3 Proofs of Theorems 1.7 and 1.9
For a bounded domain $\Omega\subset\mathbb{C}^{n}$, its Bergman kernel is defined as
$K(z,t):=\sum\varphi_{j}(z)\overline{\varphi_{j}(t)},\quad z,t\in\Omega,$
where $\{\varphi_{j}\}_{j=1}^{\infty}$ is a complete orthonormal basis for the space of $L^{2}$ holomorphic functions. The Bergman metric $B_{\Omega}(z,v)$ is defined by
$B^{2}_{\Omega}(z,v):=\sum_{\alpha,\beta=1}^{n}g_{\alpha\bar{\beta}}v_{\alpha}\overline{v_{\beta}},\quad\text{where }g_{\alpha\bar{\beta}}(z)=\frac{\partial^{2}\log K(z,z)}{\partial z_{\alpha}\partial\overline{z_{\beta}}},\quad\text{for }z\in\Omega,\ v\in\mathbb{C}^{n}.$
The Bergman metric is always a Kähler metric and it is incomplete in many cases. One can also define the Bergman kernel and metric on general complex manifolds by considering $L^{2}$ holomorphic top-forms instead of functions.
The Carathéodory metric corresponds to the analytic capacity for an open Riemann surface (see [41]).
Definition 3.1.
Let $X$ be an open hyperbolic Riemann surface. The analytic capacity of a Riemann surface $X$ is defined as
$c_{B}(z_{0})=\sup\left\{\left|{\partial f\over\partial w}(z_{0})\right|:f\in\hbox{Hol}(X,\mathbb{D}),f(z_{0})=0\right\}.$
Ahlfors introduced the analytic capacity for domains in order to study Painlevé’s question: which compact sets $E$ in the complex plane are removable for the bounded holomorphic functions?
Proof of Theorem 1.7, Part $\bf A$.
Since $\Omega$ is universally covered by a disc, in a local coordinate at $z\in X$, let $\lambda^{2}(z)|dz|^{2}$ denote the Poincaré metric with constant holomorphic sectional curvature $-2$. Due to the Ahlfors-Schwarz lemma (see [2, 37, 44]) and the fact that the Kobayashi metric dominates the Carathéodory metric, we get
$\frac{B}{2}\geq\lambda^{2}\geq c^{2}_{B},$
which yields the inequality part.
If there exists some $z_{0}\in X$ such that
$B(z_{0})=2c^{2}_{B}(z_{0}),$
then $\lambda(z_{0})=c_{B}(z_{0})$. Since the Kobayashi metric and the Carathéodory metric are equal at one point, by a lemma of the third author in [43], we conclude that $X$ is biholomorphic to a disc. The converse is straightforward as both metrics are biholomorphically invariant with explicit formulas on a disc.
We also recall the definition of the logarithmic capacity, in order to state Theorem 3.3, which is a strengthened version of Theorem 1.7, Part $\bf A$. Let $SH^{-}(X)$ denote the set of negative subharmonic functions on an open Riemann surface $X$. Given $z_{0}\in X$, let $w$ be a fixed local coordinate in a neighbourhood of $z_{0}$ such that $w(z_{0})=0$. The (negative) Green’s function is
$G(z,z_{0})=\sup\{u(z):u\in SH^{-}(X),\limsup_{z\to z_{0}}u(z)-\log|w(z)|<\infty\}.$
An open Riemann surface admits a Green’s function if and only if there is a non-constant, negative subharmonic function defined on it. The Green’s function is strictly negative on the surface and harmonic except on the diagonal. The Green’s function on a Riemann surface induces the logarithmic capacity, which is used prominently in potential theory.
Definition 3.2.
Let $X$ be an open hyperbolic Riemann surface. The logarithmic capacity $c_{\beta}$ is defined as
$c_{\beta}(z_{0})=\lim_{z\to z_{0}}\exp(G(z,z_{0})-\log|w(z)|).$
When we wish to emphasize the surface $X$, we will use the notation $c_{m;X}(\cdot)$, $m=\beta\text{ or }B$. If $h:X_{1}\to X_{2}$ is a biholomorphism, then
$c_{m;X_{1}}(z)=|h^{\prime}(z)|c_{m;X_{2}}(h(z)),\quad m=\beta\text{ or }B.$
Thus, both the logarithmic and analytic capacities are independent of the choice of the local coordinates and define conformally-invariant metrics $c_{\beta}(z)|dz|$ and $c_{B}(z)|dz|$.
Theorem 3.3.
Let $X$ be a hyperbolic Riemann surface whose Bergman metric $B(z)|dz|^{2}$ is complete with holomorphic sectional curvature no less than $-1$. In a local coordinate at $z\in X$, let
$\lambda^{2}(z)|dz|^{2},\quad c^{2}_{\beta}(z)|dz|^{2},\quad\text{and}\quad c^{2}_{B}(z)|dz|^{2}$
denote the Poincaré metric with constant holomorphic sectional curvature $-2$, the logarithmic capacity, and the analytic capacity, respectively. Then, it holds on $X$ that
$\frac{B}{2}\geq\lambda^{2}\geq c^{2}_{\beta}\geq c^{2}_{B};$ (3.1)
moreover, $X$ is biholomorphic to a disc if and only if there exists some $z_{0}\in X$ such that
$B(z_{0})=2c^{2}_{\beta}(z_{0})\quad\text{or}\quad 2c^{2}_{B}(z_{0}).$
Proof.
The first inequality in (3.1) holds due to the Ahlfors-Schwarz lemma. The second inequality is a result of Minda, who also characterized its equality part in [32], by showing that the Poincaré metric with constant holomorphic sectional curvature $-2$ coincides with the logarithmic capacity at one point if and only if the surface is simply connected. The last inequality is well known and the characterization of its equality part can be deduced from a result of Minda [32] who used the sharp form of the Lindelöf principle (see also a more direct proof of the equality characterization without adopting this principle in [15] and some related observations in [13]).
If there exists some $z_{0}\in X$ such that
$B(z_{0})=2c^{2}_{\beta}(z_{0})\quad\text{or}\quad 2c^{2}_{B}(z_{0}),$
then $\lambda(z_{0})=c_{\beta}(z_{0})$, and we conclude by [32] that $X$ is simply connected. The converse is straightforward as all the inequalities in (3.1) become equalities.
The remainder of the paper is dedicated to proving Theorem 1.7, Part $\bf B$, Corollary 1.8, and Theorem 1.9, which involve the Lu constant $L(\Omega)$.
Proof of Theorem 1.7, Part $\bf B$.
Since $\Omega$ is universally covered by covered by a ball, the Kobayashi metric on $\Omega$ is a Kähler metric with constant holomorphic sectional curvature. Due to the Ahlfors-Schwarz lemma (see [2, 37, 44]) and the lower bound of the holomorphic sectional curvature of the complete Bergman metric, we get
$L(\Omega)B_{\Omega}\geq K_{\Omega}\geq C_{\Omega}.$
If there exists some $z_{0}\in\Omega$ such that
$L(\Omega)B_{\Omega}(z_{0},v)=C_{\Omega}(z_{0},v),\quad\text{for all\,}\,v\in T_{z_{0}}(\Omega),$
then $K_{\Omega}(z_{0},v)=C_{\Omega}(z_{0},v)$. Since the Kobayashi metric and the Carathéodory metric are equal at one point, by the Wong/Stanton theorem, we conclude that $X$ is biholomorphic to a ball. The converse is straightforward as both metrics are biholomorphically invariant with explicit formulas on a ball.
Proof of Corollary 1.8.
We will prove a negated version of the corollary, namely, $\Omega$ is biholomorphic to a Euclidean ball if and only if the holomorphic sectional curvature of the Bergman metric is no less than $-\frac{2}{n+1}$ and equality is attained in the sharp Lu’s inequality at a point in $\Omega$.
The necessity holds true because for a domain biholomorphic to a Euclidean ball in $\mathbb{C}^{n}$, the holomorphic sectional curvature of the Bergman metric is identically equal to $-\frac{2}{n+1}$ and the sharp Lu’s inequality (1.1) becomes equality everywhere.
For the sufficiency, if the holomorphic sectional curvature of the Bergman metric is no less than $-\frac{2}{n+1}$, then $L(\Omega)=\frac{1}{\sqrt{n+1}}$ following from the Ahlfors-Schwarz lemma. So, the holomorphic sectional curvature is no less than $-2(L(\Omega))^{2}$. If equality is attained in the sharp Lu’s inequality at a point in $\Omega$, then we conclude by Theorem 1.7, Part $\bf B$, that $\Omega$ is biholomorphic to a ball.
Proof of Theorem 1.9.
From the upper bound of the holomorphic sectional curvature of the Bergman metric and the Ahlfors-Schwarz lemma (see [2, 37, 44]), one obtains the first inequality in the following inequality chain
$K_{\Omega}\geq L(\Omega)B_{\Omega}\geq C_{\Omega},$
where the second inequality holds true due to the definition of $L(\Omega)$. By Theorem 2.2, there exists some $\epsilon>0$ such that: if $z\in\Omega$ and $\delta_{\Omega}(z)<\epsilon$, then there is a nonempty, open set $E_{\Omega}(z)$ of directions where the Kobayashi metric $K_{\Omega}$ and the Carathéodory metric $C_{\Omega}$ coincide. One gets on $E_{\Omega}(z)$ that
$K_{\Omega}=L(\Omega)B_{\Omega}=C_{\Omega}.$
Using the argument of the third author in [42], we know that the holomorphic sectional curvature $H(z;v)$ of the Bergman metric is equal to $-2(L(\Omega))^{2}$ for any non-zero $v\in E_{\Omega}(z)$. Since the holomorphic sectional curvature function
$v\in T_{z}(\Omega)\mapsto H(z;v)\in\mathbb{R}$
is algebraic ($z$ is fixed), we obtain that the holomorphic sectional curvature $H(z;v)$ is equal to $-2(L(\Omega))^{2}$ for all non-zero vector $v\in T_{z}(\Omega)$. This is true for all $z\in\Omega$ such that $\delta_{\Omega}(z)<\epsilon$ for a sufficiently small $\epsilon$. Then, the holomorphic sectional curvature of the Bergman metric $B_{\Omega}$ is equal to $-2(L(\Omega))^{2}$ on $\Omega\setminus\Sigma$, where $\Sigma$ is defined by (2.2). Since the Bergman metric is complete on $\Omega$ whose boundary is $C^{2}$, by Theorem 2.3, the Bergman metric has constant negative holomorphic sectional curvature on $\Omega$. Then, Lu’s theorem in [31] implies that $\Omega$ is biholomorphic to a ball.
Statements and Declarations
No financial or non-financial interests that are directly or indirectly related to the work submitted for publication were reported by the authors.
Data sharing is not applicable to this article as no datasets were generated or analysed during the current study.
Funding
The research of the first author was supported by an AMS-Simons travel grant.