### v2

# Local Rigidity of the Bergman Metric and of the Kähler Carathéodory Metric

Abstract
We prove that if the Carathéodory metric on a strictly pseudoconvex domain with a smooth boundary is locally Kähler near the boundary, then the domain is biholomorphic to a ball. We also establish a local rigidity theorem for domains with Bergman metrics of constant holomorphic sectional curvature, and highlight this relationship with the Lu constant.
footnotetext: Keywords. Carathéodory metric, Bergman metric, Kähler metric, Kähler-Einstein metric, biholomorphic map, holomorphic sectional curvature, Lu constant, strictly pseudoconvex domain, rigidity theorem Mathematics Subject Classification. Primary 32F45; Secondary 32T15, 32H02
1 Introduction
A bounded pseudoconvex domain in complex Euclidean space possesses several biholomorphically invariant metrics: the Bergman metric, the differential Carathéodory metric, the unique (up to scaling) Kähler-Einstein metric, and the differential Kobayashi-Royden metric. These metrics all coincide, up to a multiplicative constant, on the unit ball $\mathbb{B}^{n}$. However, this coincidence does not generally occur for arbitrary domains, leading to the following natural question:
Question. What are the domains where the Bergman, Carathéodory, Kähler-Einstein, and Kobayashi metrics are not pairwise distinct (up to scaling)?
For a bounded domain $\Omega$ in $\mathbb{C}^{n}$, denote by $\mathbb{D}(\Omega)$ the family of holomorphic functions from $\Omega$ into $\mathbb{D}$, and denote by $\Omega(\mathbb{D})$ the family of holomorphic mappings from $\mathbb{D}$ into $\Omega$, where $\mathbb{D}$ denotes the unit disk in $\mathbb{C}$. Recall (see [48]) that:
  1. $1$
    The differential Carathéodory metric is given by
    $\begin{split}&C_{\Omega}:\Omega\times\mathbb{C}^{n}\to\mathbb{R}^{+}\cup\{0\},\\ &C_{\Omega}(z,v)=\sup\{|df(v)|:f\in\mathbb{D}(\Omega),f(z)=0\},\end{split}$
    where $df(v)$ is measured with respect to the Poincaré metric on $\mathbb{D}$;
  2. $2$
    The differential Kobayashi metric is given by
    $\begin{split}&K_{\Omega}:\Omega\times\mathbb{C}^{n}\to\mathbb{R}^{+}\cup\{0\},\\ &K_{\Omega}(z,v)=\inf\{|t|:t\text{ is a tangent vector to }\mathbb{D}\text{ at}0,\\ &\exists f\in\Omega(\mathbb{D})\text{ with }f(0)=z\text{ and }df(t)=v\},\end{split}$
    where $t$ is measured with respect to the Poincaré metric on $\mathbb{D}$.
In the seventies, the third author observed in [47] that on a complex manifold $M$, the holomorphic sectional curvature of the differential Kobayashi metric $K_{M}$ is bounded below by $-2$, and that of the differential Carathéodory metric $C_{M}$ is bounded above by $-2$. With this observation, one can immediately conclude the following characterization of the unit ball.
Theorem 1.1 ([47]).
Let $M$ be a simply-connected complete hyperbolic complex manifold with $C_{M}=K_{M}$, assumed to be a Kähler metric. Then, $M$ is biholomorphic to a ball.
A Kähler manifold is a complex manifold with a Hermitian metric whose associated 2-form is closed. Stanton proved the following result, which is more general than Theorem 1.1.
Theorem 1.2 ([45]).
Let $M$ be a complete hyperbolic complex manifold. Assume that there is a point on $M$ at which $C_{M}$ and $K_{M}$ are equal, and one of these metrics is Hermitian and of class $C^{\infty}$. Then $M$ is biholomorphic to a ball.
Gaussier and Zimmer recently established the following beautiful relationship between local curvature conditions and global structure for complete Kähler metrics (see Corollary 1.2 and its proof in [23], and a previous result of Seshadri and Verma [44] for simply connected non-positively curved Kähler manifolds), which shows that constant negative holomorphic sectional curvature on the complement of a compact set ensures the universal cover is biholomorphic to a ball.
Theorem 1.3 ([23]).
Suppose that $M$ is a Stein manifold with $\dim_{\mathbb{C}}M\geq 2$ and $g$ is a complete Kähler metric on $M$. If there exists a compact set $\Sigma\subset M$ such that $M\backslash\Sigma$ is connected and $g$ has constant negative holomorphic sectional curvature on $M\backslash\Sigma$, then the universal cover of $M$ is biholomorphic to a ball in $\mathbb{C}^{\dim_{\mathbb{C}}M}$.
In the first main result of this paper, we discover that for Bergman metrics specifically, an even stronger phenomenon under more localized conditions - the domain itself must be biholomorphic to a ball when constant curvature holds just on an open set:
Theorem 1.4.
Let $\Omega$ be a bounded domain in $\mathbb{C}^{n}$. Suppose the Bergman metric $B^{2}_{\Omega}$ is complete and its holomorphic sectional curvature is identically equal to a negative constant $-c^{2}$ on an open set $U\subset\Omega$. Then, $\Omega$ is biholomorphic to $\mathbb{B}^{n}$, where $n=2c^{-2}-1$.
This observation advances the classical work in this area. A famous theorem of Lu [36] says that a bounded domain in $\mathbb{C}^{n}$ with a complete Bergman metric of constant holomorphic sectional curvature must be biholomorphic to $\mathbb{B}^{n}$. While the global constant curvature follows from the identity theorem of real-analytic functions, our proof provides an explicit construction of the Bergman representative coordinate $T$ as a holomorphic isometry, which has independent interest and applications. Similar uniqueness principles based on real-analyticity appear in our analysis of strictly pseudoconvex domains. Our result shows that Lu’s conclusion holds under weaker hypotheses, requiring the constant curvature condition only on an open set. In [18, 19], the first and third authors extended Lu’s theorem to the Bergman-incomplete situation and provided a multidimensional generalization of Carathéodory’s theorem on the continuous extension of biholomorphisms up to the closures. See also the recent works [28] by Huang and Li, and [21] by Ebenfelt, Treuer and Xiao.
These geometric insights naturally lead us to examine the quantitative relationships between different metrics on bounded domains. A foundational result in this direction is Lu’s inequality, which establishes precise control of the Carathéodory metric by the Bergman metric:
Theorem 1.5 ([35], cf. [10]).
Let $\Omega$ be a bounded domain in $\mathbb{C}^{n}$. For each tangent vector $v\in T_{z}(\Omega)=\mathbb{C}^{n}$ at $z\in\Omega$, it holds that
$B_{\Omega}(z,v)\geq C_{\Omega}(z,v),$
where $B_{\Omega}(z,v)$ and $C_{\Omega}(z,v)$ are the lengths of $v$ with respect to the Bergman metric $B^{2}_{\Omega}$ and the differential Carathéodory metric $C_{\Omega}$, respectively.
The above inequality due to Lu, while fundamental, is not sharp. To obtain optimal control of the relationship between these metrics, we consider the following Lu constant.
Definition. The Lu constant $L(\Omega)$ of a bounded domain $\Omega$ in $\mathbb{C}^{n}$ is defined as
[…]
The definition of $L(\Omega)$ implies that
$L(\Omega)B_{\Omega}\geq C_{\Omega}.$ (1.1)
For the rest of this paper, we shall call (1.1) the sharp Lu’s inequality. The value of the Lu constant $L(\Omega)$ is understood in the following scenarios:
By Theorem 1.5, $L(\Omega)\leq 1$. Also, $L(\Omega)=\frac{1}{\sqrt{n+1}}$ when $\Omega$ is the unit ball, and $L(\Omega)=\sqrt{2n}$ when $\Omega$ is the unit polydisc $\mathbb{D}^{n}$. One further determines the Lu constants of all Cartan classical domains. When $\Omega$ is a bounded strictly pseudoconvex domain in $\mathbb{C}^{n}$, $L(\Omega)\geq\frac{1}{\sqrt{n+1}}$ by the results on the boundary behaviors of $B_{\Omega}$ and $C_{\Omega}$ due to Diederich [20] and Graham [24], respectively. If the holomorphic sectional curvature of the Bergman metric on a bounded strictly pseudoconvex domain in $\mathbb{C}^{n}$ is bounded below by $-\frac{2}{n+1}$, then $L(\Omega)=\frac{1}{\sqrt{n+1}}$ following from the Ahlfors-Schwarz lemma.
Next, along the line of the Wong/Stanton theorem, we present a rigidity result on the comparison between the Bergman metric and the Carathéodory metric.
Theorem 1.6.
Let $X$ be a hyperbolic Riemann surface. Assume that the Bergman metric $B(z)|dz|^{2}$ is complete with holomorphic sectional curvature no less than $-1$, and the Carathéodory metric is written as $c^{2}_{B}(z)|dz|^{2}$, where $z$ is a local coordinate. Then, it holds on $X$ that
${B}\geq 2c^{2}_{B};$
moreover, $X$ is biholomorphic to a disc if and only if there exists some $z_{0}\in X$ such that
$B(z_{0})=2c^{2}_{B}(z_{0}).$
Let $\Omega\subset\mathbb{C}^{n}$ be a domain whose universal cover is biholomorphic to a Euclidean ball. Assume that the Bergman metric $B^{2}_{\Omega}$ is complete on $\Omega$ with holomorphic sectional curvature no less than $-2(L(\Omega))^{2}$. Then, $\Omega$ is biholomorphic to a ball if and only if there exists some $z_{0}\in\Omega$ such that
$L(\Omega)B_{\Omega}(z_{0},v)=C_{\Omega}(z_{0},v),\quad\text{for all\,}\,v\in T_{z_{0}}(\Omega).$
Since a bounded pseudoconvex domain with a spherical boundary in $\mathbb{C}^{n}$, $n\geq 2$, is universally covered by a ball by the uniformization theorem (see [41, Theorem A.2]), we obtain the following corollary relating to the examples of Burns and Shnider [8] using the aforementioned methods:
Corollary 1.7.
Let $\Omega$ be a bounded strictly pseudoconvex domain with spherical boundary in $\mathbb{C}^{n}$, $n\geq 2$. Then $\Omega$ is not biholomorphic to a Euclidean ball if and only if either the holomorphic sectional curvature of the Bergman metric in some direction is strictly less than $-\frac{2}{n+1}$ or the sharp Lu’s inequality (1.1) is strict at all points and for some directions.
In the second half of the paper, we turn our attention to bounded domains where the Carathéodory metric exhibits special geometric properties - specifically, when it coincides up to scaling with the Bergman or Kähler-Einstein metric, or more generally, when it is a Kähler metric. This investigation connects to a broader question raised (in a slightly different form) by Yau [50, pp. 679] concerning the characterization of manifolds whose Bergman metrics are Kähler-Einstein.
On strictly pseudoconvex domains, significant progress has been made in understanding the relationships between these metrics. Cheng and Yau [11] showed the existence of complete Kähler-Einstein metrics on smoothly bounded strictly pseudoconvex domains in $\mathbb{C}^{n}$, while Mok and Yau [38] later extended this to general bounded pseudoconvex domains without boundary regularity conditions. For smoothly bounded strictly pseudoconvex domains, Cheng conjectured that the Bergman metric is Kähler-Einstein if and only if the domain is biholomorphic to a ball. Following the earlier works of Fu and the third author [22] and Nemirovski and Shafikov [41], this conjecture was confirmed by Huang and Xiao in [29]. Additionally, see the recent work [43] by Savale and Xiao for finite type domains in $\mathbb{C}^{2}$.
Building on these foundations and inspired by recent work of Gaussier and Zimmer [23], who characterized domains whose universal coverings are biholomorphic to a ball through properties of the Kobayashi metric, we establish parallel results for the Carathéodory metric. Their characterization showed:
Theorem 1.8 ([23]).
Suppose that $\Omega\subset\mathbb{C}^{n}$, $n\geq 2$, is a bounded strictly pseudoconvex domain with a $C^{2}$ boundary. Then, the following are equivalent:
  1. $(a)$
    The Kobayashi metric on $\Omega$ is a Kähler metric;
  2. $(b)$
    The Kobayashi metric on $\Omega$ is a Kähler metric with constant holomorphic sectional curvature;
  3. $(c)$
    The universal cover of $\Omega$ is biholomorphic to a ball;
  4. $(d)$
    The Kobayashi metric on $\Omega$ is a Kähler-Einstein metric.
Moreover, the Kobayashi metric on $\Omega$ is a scalar multiple of the Bergman metric if and only if $\Omega$ is biholomorphic to a ball.
For strictly pseudoconvex domains, we obtain the following main result, which is a Carathéodory-analogue of both Gaussier-Zimmer’s theorem and the resolution of the Cheng conjecture:
Theorem 1.9.
Suppose that $\Omega\subset\mathbb{C}^{n}$, $n\geq 2$, is a bounded strictly pseudoconvex domain with a $C^{2}$ boundary. Then, the following are equivalent:
  1. $(1)$
    The Carathéodory metric on $\Omega$ is a Kähler metric on $U\cap\Omega$, where $U$ is an open set that contains the boundary of $\Omega$;
  2. $(2)$
    The Carathéodory metric on $\Omega$ is a Kähler metric with constant holomorphic sectional curvature;
  3. $(3)$
    $\Omega$ is biholomorphic to the unit ball;
  4. $(4)$
    The Carathéodory metric on $\Omega$ is a scalar multiple of the Bergman metric;
  5. $(5)$
    The Carathéodory metric on $\Omega$ is a Kähler-Einstein metric.
For $n=1$, Suita [46] (cf. [6]) showed that the Carathéodory metric is real analytic for any bounded domain in $\mathbb{C}$. Thus, it can be easily proven that if the Carathéodory metric is complete with holomorphic sectional curvature identically equal to $-2$, then the domain is biholomorphic to a disc.
Some related examples are discussed below.
Examples.
  1. $(i)$
    By the remarkable works of Lempert [33, 31, 34], the Carathéodory metric and the Kobayashi metric coincide on convex domains. Theorem 1.1/1.2 implies that the Carathéodory metric on a bounded convex domain is a Hermitian metric if and only if the domain is biholomorphic to the unit ball. In particular, let $\mathcal{P}$ be a small generic perturbation of $\mathbb{B}^{n}$ such that $\mathcal{P}$ is convex in $\mathbb{C}^{n}$ but not biholomorphic to $\mathbb{B}^{n}$ when $n\geq 2$ (see the papers of Burns, Shnider and Wells [9], and Greene and Krantz[26]). Then, the Carathéodory metric on $\mathcal{P}$ is not a Hermitian metric when $n\geq 2$.
    In [16], the first author and Treuer constructed a bounded complete Reinhardt domain
    :=$\mathcal{R}:=\{z\in\mathbb{C}^{2}:|z_{1}|^{4}+|z_{1}|^{2}+|z_{2}|^{2}<1\}$
    which is strongly convex with an algebraic boundary and not biholomorphic to $\mathbb{B}^{2}$. By the above discussion, the Carathéodory metric on $\mathcal{R}$ is not a Hermitian metric.
    Also, note that a complex analytic ellipsoid
    ${E}:=\{(z,w)\in{\mathbb{C}^{1}}\times{\mathbb{C}^{1}}:|z|^{2}+|w|^{2p}<1\},$
    is convex and not biholomorphic to $\mathbb{B}^{2}$, where $\mathbb{Z}^{+}\ni p>1$. So, the Carathéodory metric on ${E}$ is not a Hermitian metric. (See some recent related work in this direction by Cho [14]).
  2. $(ii)$
    The Kobayashi metric of the unit bidisc is
    $K_{\mathbb{D}^{2}}((z,w),(X,Y))=\max\{K_{\mathbb{D}}(z,X),K_{\mathbb{D}}(w,Y)\}.$
    Thus, the Carathéodory metric and the Kobayashi metric are not $C^{1}$-smooth metrics in general. For bounded symmetric domains of rank $\geq 2$, the Carathéodory and Kobayashi metrics are not Kähler.
  3. $(iii)$
    Burns and Shnider [8] have constructed examples of non-simply connected strictly pseudoconvex domains with spherical boundaries in $\mathbb{C}^{n}$. By the uniformization theorem (see [39, 40]), such domains are universally covered by the unit ball, so the Kobayashi metric is Kähler-Einstein. For example, let
    $\mathcal{S}:=\{(z,w)\in\mathbb{C}^{2}\mid\sin\log|z|+|w|^{2}<0,e^{-\pi}<|z|<1\}.$
    By our Theorem 1.9 above, the Carathéodory metric $C_{\mathcal{S}}$ is not a Kähler metric on $\mathcal{S}$, and thus $C_{\mathcal{S}}$ is not equal to the Kobayashi metric $K_{\mathcal{S}}$.
Lastly, we extend a previous result [13] (see also [12]) by Cheung and the third author from bounded convex domains to strictly pseudoconvex domains.
Theorem 1.10.
Let $\Omega\subset\mathbb{C}^{n}$, $n\geq 2$, be a bounded strictly pseudoconvex domain with a $C^{2}$ boundary and the Lu constant $L(\Omega)$. Suppose the holomorphic sectional curvature of the Bergman metric $B^{2}_{\Omega}$ on $\Omega$ is bounded above by $-2(L(\Omega))^{2}$. Then $\Omega$ is biholomorphic to a ball.
Our proofs of Theorems 1.9 and 1.10 depend heavily on the work of Bracci, Fornæss, and Wold [5] and that of Gaussier and Zimmer [23]. (For related observations, see earlier works by Lempert [32] and Huang [27].) One of the key ideas is that at a point sufficiently close to the boundary of a smoothly bounded strictly pseudoconvex domain, there is a nonempty, open set of directions where the Kobayashi and Carathéodory metrics coincide. We observe that the Carathéodory metric has constant negative holomorphic sectional curvature in these directions, and in fact, has constant negative holomorphic sectional curvature near the boundary, if the metric is Kähler. The proof of Theorem 1.10 also relies on our Theorem 1.4.
The organization of the paper is as follows. In Section 2, we give a proof of our main result (Theorem 1.9) characterizing when the Carathéodory metric is Kähler. In Section 3, we establish our local-to-global principle for the Bergman metric (Theorem 3.4) which yields our first main theorem (Theorem 1.4), and then provide proofs of our rigidity results (Theorems 1.6 and 1.10) relating the Lu constant to geometric properties of domains.
2 When is the Carathéodory metric a Kähler metric?
[…]
$\delta_{\Omega}(z)=\min\{\|z-x\|:x\in\partial\Omega\}.$
Definition 2.1.
A holomorphic map $\varphi:\mathbb{D}\rightarrow\Omega$ is a complex geodesic with respect to the Kobayashi (or Carathéodory) metric if
$k_{\Omega}(\varphi(z_{1}),\varphi(z_{2}))=k_{\mathbb{D}}(z_{1},z_{2})\quad(\text{or \,}c_{\Omega}(\varphi(z_{1}),\varphi(z_{2}))=c_{\mathbb{D}}(z_{1},z_{2}))$
for all $z_{1},z_{2}\in\mathbb{D}$, where $k_{\Omega}$ (or $c_{\Omega}$) denotes the Kobayashi (or Carathéodory) distance.
Suppose that ${C_{\Omega}(z,v)}={K_{\Omega}(z,v)}$ for $z\in\Omega,\,v\in T_{z}(\Omega)$, where $\Omega$ is a bounded domain in $\mathbb{C}^{n}$. By the definitions of the Carathéodory and Kobayashi metrics, let $\rho:\Omega\to\mathbb{D}$ be a holomorphic function that realizes the Carathéodory metric $C_{\Omega}(z,v)$, and let $\varphi:\mathbb{D}\to\Omega$ be a holomorphic map that realizes the Kobayashi metric $K_{\Omega}(z,v)$. Then, the composition $\rho\circ\varphi$ is an automorphism of $\mathbb{D}$ that fixes the origin. It can be shown that $\varphi:\mathbb{D}\to\Omega$ is a complex geodesic with respect to both the Carathéodory and Kobayashi metrics. Furthermore, the holomorphic sectional curvatures of $C_{\Omega}$ and $K_{\Omega}$ restricted to the image $\varphi(\mathbb{D})$ are both equal to $-2$ everywhere.
When $\Omega$ has a $C^{2}$ boundary and $z\in\overline{\Omega}$ is sufficiently close to $\partial\Omega$, there is a unique point $\pi(z)\in\partial\Omega$ such that $\|z-\pi(z)\|=\delta_{\Omega}(z)$. We let
$P_{z}:\mathbb{C}^{n}\rightarrow T_{\pi(z)}^{\mathbb{C}}\partial\Omega$
denote the (Euclidean) orthogonal projection and let $P_{z}^{\bot}=\text{Id}-P_{z}$. The following result states that complex geodesics exist at base points sufficiently close to the smooth boundary of strictly pseudoconvex domains and in directions sufficiently tangential.
Theorem 2.2.
Suppose that $\Omega\subset\mathbb{C}^{n},n\geq 2,$ is a bounded strictly pseudoconvex domain with $C^{2}$ boundary. Then there exists $\epsilon>0$ such that: if $z\in\Omega$ and $\delta_{\Omega}(z)<\epsilon$, then there exists an non-empty open set $E_{\Omega}(z)$ of non-zero tangent vectors at $z$ where the Kobayashi and Carathéodory metrics agree.
Theorem 2.2 is fundamental for our analysis. The existence of complex geodesics in strictly pseudoconvex domains was first established by Lempert [32], who showed that the relevant estimates depend only on the normal curvature of the domain. Lempert’s original analysis demonstrated that since all estimates depend only on normal curvature, the results naturally extend to the $C^{2}$ case through Montel’s theorem, even though his original proof was given for the analytic case. Moreover, Huang [27] proved that certain extremal mappings near boundary points are complex geodesics (denoted here as $\varphi$). See also [7, 3] for related results. The existence of an open set where the Kobayashi and Carathéodory metrics coincide was first shown by Kosiński [30]. Later for $C^{3}$ domains, Bracci, Fornæss, and Wold [5, Theorem 1.1] established the existence of the left inverse (denoted here as $\rho$) and provided precise conditions for when the Kobayashi and Carathéodory metrics coincide near the boundary by showing
${C_{\Omega}(z,v)}={K_{\Omega}(z,v)},\quad z\in\Omega,\,v\in T_{z}(\Omega)$
if $\delta_{\Omega}(z)<\epsilon$ and if
$\left\lVert P_{z}^{\bot}(v)\right\rVert<\epsilon\left\lVert P_{z}(v)\right\rVert.$
The proof of Theorem 2.2 was provided by Gaussier and Zimmer [23], who also obtained the following result which will play a crucial role in our arguments.
Theorem 2.3 ([23]).
Suppose that $M$ is a Stein manifold with $\dim_{\mathbb{C}}M\geq 2$, $\Sigma\subset M$ is a compact set where $M\backslash\Sigma$ is connected, and $g_{0}$ is a Hermitian metric on $M\backslash\Sigma$ which is complete at infinity. If $g_{0}$ is locally symmetric metric, then there exists a complete locally symmetric Hermitian metric $g$ on $M$ such that $g=g_{0}$ on $M\backslash\Sigma$.
Theorem 2.3 implies Theorem 1.3 since a Kähler metric with constant holomorphic sectional curvature is locally symmetric.
Proof of Theorem 1.9.
Since $(2)$ implies $(1)$ by definition, and it is well known that $(3)$ implies $(2)$, we only need to prove the non-trivial part, namely that $(1)$ implies $(3)$. By Theorem 2.2, there exists some $\epsilon>0$ such that: if $z\in\Omega$ and $\delta_{\Omega}(z)<\epsilon$, then there is a nonempty, open set $E_{\Omega}(z)$ of directions where the Carathéodory metric and the Kobayashi metric coincide. Using the argument in [47], we know that the holomorphic sectional curvature $H(z;v)$ of the Carathéodory metric is equal to $-2$ for any non-zero $v\in E_{\Omega}(z)$. Briefly speaking, given that the Carathéodory metric and Kobayashi metric coincide at $z$ for $v$, there exists a complex geodesic with respect to the Carathéodory metric on which the holomorphic sectional curvature is equal to $-2$ at every point. As the curvature $H(z;v)$ of $\Omega$ is bounded above by $-2$, the curvature-decreasing property implies that it must be exactly $-2$ at $z$. Moreover, since the holomorphic sectional curvature function
$v\in T_{z}(\Omega)\mapsto H(z;v)\in\mathbb{R}$
is algebraic (with $z$ fixed), it follows that the holomorphic sectional curvature $H(z;v)$ equals $-2$ for all non-zero vectors $v\in T_{z}(\Omega)$. This is true for all $z\in\Omega$ such that $\delta_{\Omega}(z)<\epsilon$ for a sufficiently small $\epsilon$. Define the compact set
$\Sigma:=\{z\in\Omega:\delta_{\Omega}(z)\geq\epsilon\}$ (2.1)
as in Theorem 1.3. Then, the holomorphic sectional curvature of the Carathéodory metric $C_{\Omega}$ is equal to $-2$ on $\Omega\setminus\Sigma$.
By the work of Graham [24, Proposition 5], the Carathéodory metric $C_{\Omega}$ is complete on $\Omega$. If $C_{\Omega}$ is a Kähler metric on $U\cap\Omega$, then by Theorem 2.3, there exists a complete locally symmetric metric ${g}$ on $\Omega$ such that such that
${g}=C_{\Omega}\quad\text{on}\quad U\cap\Omega.$ (2.2)
Since the universal cover of $(\Omega,{g})$ is globally symmetric (hence homogeneous) and ${g}$ has constant negative holomorphic sectional curvature on $U\cap\Omega$, we then see that ${g}$ has constant negative holomorphic sectional curvature on $\Omega$. Thus, the universal cover of $\Omega$ is biholomorphic to the unit ball. By Theorem 1.8, the Kobayashi metric $K_{\Omega}$ is a complete Kähler metric of constant negative holomorphic sectional curvature on $\Omega$. Let $\pi:\mathbb{B}^{n}\rightarrow\Omega$ be the holomorphic covering map, which is a local isometry for $K_{\Omega}$. Hence, we have $\pi^{*}K_{\Omega}=\pi^{*}{g}$ with holomorphic sectional curvature equaling $-2$ on $\mathbb{B}^{n}$. At any point $z\in\Omega\setminus\Sigma$, it holds by (2.2) that
${C_{\Omega}(z,v)}={{g}(z,v)}={K_{\Omega}(z,v)},\quad v\in T_{z}(\Omega).$
By Theorem 1.2, one concludes that $\Omega$ is biholomorphic to a ball.
Regarding Conditions $(4)$ and $(5)$, notice that Condition $(3)$ implies both of them. And either Condition $(4)$ or Condition $(5)$ would imply Condition $(1)$ by definition. Thus, the proof is complete.
From the proof of Theorem 1.9, we see that Condition $(a)$ in Theorem 1.8 can be relaxed to:
$(a^{\prime})$ The Kobayashi metric on $\Omega$ is a Kähler metric on $U\cap\Omega$, where $U$ is an open set that contains the boundary of $\Omega$.
Remark. On April 19th, 2024, the first author gave an online talk at Syracuse University’s Analysis Seminar hosted by Yuan Yuan. During the talk, Yuan mentioned that he had obtained some related result yet to be written up. The results in this paper were obtained around 2022/2023.
3 Local-to-global principle for the Bergman metric
For a bounded domain $\Omega\subset\mathbb{C}^{n}$, its Bergman kernel is defined as
$K(z,t):=\sum\varphi_{j}(z)\overline{\varphi_{j}(t)},\quad z,t\in\Omega,$
where $\{\varphi_{j}\}_{j=1}^{\infty}$ is a complete orthonormal basis for the space of $L^{2}$ holomorphic functions. The Bergman metric $B^{2}_{\Omega}(z,v)$ is defined by
$B^{2}_{\Omega}(z,v):=\sum_{\alpha,\beta=1}^{n}g_{\alpha\bar{\beta}}v_{\alpha}\overline{v_{\beta}},\quad\text{where }g_{\alpha\bar{\beta}}(z)=\frac{\partial^{2}\log K(z,z)}{\partial z_{\alpha}\partial\overline{z_{\beta}}},\quad\text{for }z\in\Omega,\ v\in\mathbb{C}^{n}.$
The Bergman metric is always a Kähler metric and it is incomplete in many cases. One can also define the Bergman kernel and metric on general complex manifolds by considering $L^{2}$ holomorphic top-forms instead of functions. The Bergman representative coordinate $T(z)=(w_{1},...,w_{n})^{\tau}$ relative to $p\in\Omega$ is defined as
$w_{\alpha}(z):=\sum_{j=1}^{n}g^{\bar{j}\alpha}(p)\left(K(z,p)^{-1}\left.\frac{\partial}{\partial\overline{t_{j}}}\right|_{t=p}K(z,t)-\left.\frac{\partial}{\partial\overline{t_{j}}}\right|_{t=p}\log K(t,t)\right),$ (3.1)
where $(g^{\bar{j}\alpha})=(g_{\alpha\bar{j}})^{-1}$. It is well known that $T(z)$ is holomorphic on $\Omega$ less the zero set of $K(\cdot,p)$, cf. [25].
The Carathéodory metric corresponds to the analytic capacity for an open Riemann surface (see [46]).
Definition 3.1.
Let $X$ be an open hyperbolic Riemann surface. The analytic capacity of a Riemann surface $X$ is defined as
$c_{B}(z_{0})=\sup\left\{\left|{\partial f\over\partial w}(z_{0})\right|:f\in\hbox{Hol}(X,\mathbb{D}),f(z_{0})=0\right\}.$
Ahlfors introduced the analytic capacity for domains in order to study Painlevé’s question: which compact sets $E$ in the complex plane are removable for the bounded holomorphic functions?
Proof of Theorem 1.6, Part $\bf A$.
Since $\Omega$ is universally covered by a disc, in a local coordinate at $z\in X$, let $\lambda^{2}(z)|dz|^{2}$ denote the Poincaré metric with constant holomorphic sectional curvature $-2$. Due to the Ahlfors-Schwarz lemma (see [2, 42, 49]) and the fact that the Kobayashi metric dominates the Carathéodory metric, we get
$\frac{B}{2}\geq\lambda^{2}\geq c^{2}_{B},$
which yields the inequality part.
If there exists some $z_{0}\in X$ such that
$B(z_{0})=2c^{2}_{B}(z_{0}),$
then $\lambda(z_{0})=c_{B}(z_{0})$. Since the Kobayashi metric and the Carathéodory metric are equal at one point, by a lemma of the third author in [48], we conclude that $X$ is biholomorphic to a disc. The converse is straightforward as both metrics are biholomorphically invariant with explicit formulas on a disc.
We also recall the definition of the logarithmic capacity, in order to state Theorem 3.3, which is a strengthened version of Theorem 1.6, Part $\bf A$. Let $SH^{-}(X)$ denote the set of negative subharmonic functions on an open Riemann surface $X$. Given $z_{0}\in X$, let $w$ be a fixed local coordinate in a neighbourhood of $z_{0}$ such that $w(z_{0})=0$. The (negative) Green’s function is
$G(z,z_{0})=\sup\{u(z):u\in SH^{-}(X),\limsup_{z\to z_{0}}u(z)-\log|w(z)|<\infty\}.$
An open Riemann surface admits a Green’s function if and only if there is a non-constant, negative subharmonic function defined on it. The Green’s function is strictly negative on the surface and harmonic except on the diagonal. The Green’s function on a Riemann surface induces the logarithmic capacity, which is used prominently in potential theory.
Definition 3.2.
Let $X$ be an open hyperbolic Riemann surface. The logarithmic capacity $c_{\beta}$ is defined as
$c_{\beta}(z_{0})=\lim_{z\to z_{0}}\exp(G(z,z_{0})-\log|w(z)|).$
When we wish to emphasize the surface $X$, we will use the notation $c_{m;X}(\cdot)$, $m=\beta\text{ or }B$. If $h:X_{1}\to X_{2}$ is a biholomorphism, then
$c_{m;X_{1}}(z)=|h^{\prime}(z)|c_{m;X_{2}}(h(z)),\quad m=\beta\text{ or }B.$
Thus, both the logarithmic and analytic capacities are independent of the choice of the local coordinates and define conformally-invariant metrics $c_{\beta}(z)|dz|$ and $c_{B}(z)|dz|$.
Theorem 3.3.
Let $X$ be a hyperbolic Riemann surface whose Bergman metric $B(z)|dz|^{2}$ is complete with holomorphic sectional curvature no less than $-1$. In a local coordinate at $z\in X$, let
$\lambda^{2}(z)|dz|^{2},\quad c^{2}_{\beta}(z)|dz|^{2},\quad\text{and}\quad c^{2}_{B}(z)|dz|^{2}$
denote the Poincaré metric with constant holomorphic sectional curvature $-2$, the logarithmic capacity, and the analytic capacity, respectively. Then, it holds on $X$ that
$\frac{B}{2}\geq\lambda^{2}\geq c^{2}_{\beta}\geq c^{2}_{B};$ (3.2)
moreover, $X$ is biholomorphic to a disc if and only if there exists some $z_{0}\in X$ such that
$B(z_{0})=2c^{2}_{\beta}(z_{0})\quad\text{or}\quad 2c^{2}_{B}(z_{0}).$
Proof.
The first inequality in (3.2) holds due to the Ahlfors-Schwarz lemma. The second inequality is a result of Minda, who also characterized its equality part in [37], by showing that the Poincaré metric with constant holomorphic sectional curvature $-2$ coincides with the logarithmic capacity at one point if and only if the surface is simply connected. The last inequality is well known and the characterization of its equality part can be deduced from a result of Minda [37] who used the sharp form of the Lindelöf principle (see also a more direct proof of the equality characterization without adopting this principle in [17] and some related observations in [15]).
If there exists some $z_{0}\in X$ such that
$B(z_{0})=2c^{2}_{\beta}(z_{0})\quad\text{or}\quad 2c^{2}_{B}(z_{0}),$
then $\lambda(z_{0})=c_{\beta}(z_{0})$, and we conclude by [37] that $X$ is simply connected. The converse is straightforward as all the inequalities in (3.2) become equalities.
Next, we prove Theorem 1.6, Part $\bf B$ and Corollary 1.7, which involve the Lu constant $L(\Omega)$.
Proof of Theorem 1.6, Part $\bf B$.
Since $\Omega$ is holomorphically covered by a ball, the Kobayashi metric on $\Omega$ is a Kähler metric with constant holomorphic sectional curvature. Due to the Ahlfors-Schwarz lemma (see [2, 42, 49]) and the lower bound of the holomorphic sectional curvature of the complete Bergman metric, we get
$L(\Omega)B_{\Omega}\geq K_{\Omega}\geq C_{\Omega}.$
If there exists some $z_{0}\in\Omega$ such that
$L(\Omega)B_{\Omega}(z_{0},v)=C_{\Omega}(z_{0},v),\quad\text{for all\,}\,v\in T_{z_{0}}(\Omega),$
then $K_{\Omega}(z_{0},v)=C_{\Omega}(z_{0},v)$. Since the Kobayashi metric and the Carathéodory metric are equal at one point, by the Wong/Stanton theorem, we conclude that $X$ is biholomorphic to a ball. The converse is straightforward as both metrics are biholomorphically invariant with explicit formulas on a ball.
Proof of Corollary 1.7.
We will prove a negated version of the corollary, namely, $\Omega$ is biholomorphic to a Euclidean ball if and only if the holomorphic sectional curvature of the Bergman metric is no less than $-\frac{2}{n+1}$ and equality is attained in the sharp Lu’s inequality at a point in $\Omega$.
The necessity holds true because for a domain biholomorphic to a Euclidean ball in $\mathbb{C}^{n}$, the holomorphic sectional curvature of the Bergman metric is identically equal to $-\frac{2}{n+1}$ and the sharp Lu’s inequality (1.1) becomes equality everywhere.
For the sufficiency, if the holomorphic sectional curvature of the Bergman metric is no less than $-\frac{2}{n+1}$, then $L(\Omega)=\frac{1}{\sqrt{n+1}}$ following from the Ahlfors-Schwarz lemma. So, the holomorphic sectional curvature is no less than $-2(L(\Omega))^{2}$. If equality is attained in the sharp Lu’s inequality at a point in $\Omega$, then we conclude by Theorem 1.6, Part $\bf B$, that $\Omega$ is biholomorphic to a ball.
The remainder of the paper is dedicated to understanding domains where the Bergman metric has constant negative holomorphic sectional curvature on an open set. We first establish a fundamental result about such domains through a local-to-global principle (Theorem 3.4), which leads to our characterization theorem (Theorem 1.4).
Theorem 3.4 (Local-to-global principle for Bergman metrics).
Let $\Omega$ be a bounded domain in $\mathbb{C}^{n}$. Suppose the holomorphic sectional curvature of the Bergman metric $B^{2}_{\Omega}$ is identically equal to a negative constant $-c^{2}$ on an open set $U\subset\Omega$. Then, for any $p\in U$, the Bergman representative coordinate $T(z)=(w_{1},...,w_{n})^{\tau}$ defined by (3.1) is a bounded holomorphic map from $\Omega$ to the ball
:=$\mathcal{B}:=\{(w_{1},...,w_{n})^{\tau}:\sum_{\alpha,\beta=1}^{n}w_{\alpha}g_{\alpha\bar{\beta}}(p)\overline{w_{\beta}}<{2}{c^{-2}}\}.$
Moreover, $T:(\Omega,B^{2}_{\Omega})\to(\mathcal{B},\omega_{\mathcal{B}})$ is a holomorphic isometry, where $\omega_{\mathcal{B}}$ is the Bergman metric of $\mathcal{B}$. Consequently, the holomorphic sectional curvature of the Bergman metric $B^{2}_{\Omega}$ is identically equal to a negative constant $-c^{2}$ on $\Omega$.
Proof.
First, we assume that the Bergman metric at $p\in U$ satisfies
$g_{\alpha\bar{\beta}}(p)=\delta_{\alpha\beta},$ (3.3)
and in this case we will show that $T(z)$ defined by (3.1) is a bounded holomorphic map from $\Omega$ to the ball $\mathbb{B}^{n}:=\{w\in\mathbb{C}^{n}:|w|^{2}<{2}{c^{-2}}\}$.
At any $p\in U$, by the uniqueness of the Taylor expansion (see Bochner [4] and Lu [36, Lemma 3]), there exists a smaller neighbourhood $U_{p}$ such that the Bergman kernel can be locally decomposed as
$K(z,z)=\left(1-\frac{c^{2}}{2}|T(z)|^{2}\right)^{\frac{-2}{c^{2}}}e^{f(T(z))+\overline{f(T(z))}},\quad z\in U_{p},$ (3.4)
where $f$ is holomorphic on $U_{p}$. The map $T$ defined by (3.1) is holomorphic on $\Omega\setminus A_{p}$, where $A_{p}:=\{z\in\Omega\,|\,K(z,p)=0\,\}$ is the zero set of the Bergman kernel $K(\cdot,p)$. Let $\Omega^{\prime}:=\{z\in\Omega\setminus A_{p}:T(z)\in\mathbb{B}^{n}\}$ be the set of points in $\Omega\setminus A_{p}$ that are mapped into the ball. In particular, $U_{p}\subset\Omega^{\prime}$. By (3.4) and the theory of power series, one may duplicate the variable with its conjugate so that the full Bergman kernel can be complex analytically continued as
$K(z,{z_{0}})=\left(1-\frac{c^{2}}{2}\sum_{\alpha=1}^{n}w_{\alpha}(z)\overline{w_{\alpha}(z_{0})}\right)^{\frac{-2}{c^{2}}}e^{f(T(z))+\overline{f(T({z_{0}}))}},\quad z,z_{0}\in U_{p}.$
Consider two Kähler potentials
$\Phi_{p}(z):=\log\frac{K(z,z)K(p,p)}{|K(z,p)|^{2}},\quad z\in\Omega\setminus A_{p}$
and
$\Phi_{z_{0}}(z):=\log\frac{K(z,z)K({z_{0}},{z_{0}})}{|K(z,{z_{0}})|^{2}},\quad z\in\Omega\setminus A_{z_{0}}$
for the Bergman metric $B^{2}_{\Omega}=\partial\overline{\partial}\Phi_{z_{0}}=\partial\overline{\partial}\Phi_{p}$, for $z\in(\Omega\setminus A_{p})\cap(\Omega\setminus A_{z_{0}})$. When $U_{p}$ is small enough, for any $z_{0}\in U_{p}$, it holds that
$\Phi_{z_{0}}(z)$ $=\log\frac{\left(1-\frac{c^{2}}{2}|T(z)|^{2}\right)^{\frac{-2}{c^{2}}}e^{f(T(z))+\overline{f(T(z))}}\left(1-\frac{c^{2}}{2}|T(z_{0})|^{2}\right)^{\frac{-2}{c^{2}}}e^{f(T(z_{0}))+\overline{f(T(z_{0}))}}}{\left|1-\frac{c^{2}}{2}\sum_{\alpha=1}^{n}w_{\alpha}(z)\overline{w_{\alpha}(z_{0})}\right|^{\frac{-4}{c^{2}}}|e^{f(T(z))+\overline{f(T({z_{0}}))}}|^{2}}$
$={\frac{-2}{c^{2}}}\log\left[\left(1-\frac{c^{2}}{2}|T(z)|^{2}\right)\left(1-\frac{c^{2}}{2}|T(z_{0})|^{2}\right)\left|1-\frac{c^{2}}{2}\sum_{\alpha=1}^{n}w_{\alpha}(z)\overline{w_{\alpha}(z_{0})}\right|^{-2}\right],\quad z\in U_{p},$
which yields that
$\Phi_{p}(z_{0})=\Phi_{z_{0}}(p)={\frac{-2}{c^{2}}}\log\left(1-\frac{c^{2}}{2}|T(z_{0})|^{2}\right).$
Note that $\Phi_{p}(z)$ is defined on $\Omega\setminus A_{p}$ and thus on $\Omega^{\prime}$, where ${\frac{-2}{c^{2}}}\log\left(1-\frac{c^{2}}{2}|T(z)|^{2}\right)$ can be defined. Since these two real-analytic functions coincide on $U_{p}$, by the identity theorem they are identical to each other on $\Omega^{\prime}$. That is,
$\Phi_{p}(z)={\frac{-2}{c^{2}}}\log\left(1-\frac{c^{2}}{2}|T(z)|^{2}\right),\quad z\in\Omega^{\prime}.$ (3.5)
1) We claim that no point in $\Omega\setminus A_{p}$ is mapped outside the ball $\mathbb{B}^{n}$ by $T$.
If not, suppose there exists some point $q\in\Omega\setminus A_{p}$ that is mapped to $\{w\in\mathbb{C}^{n}:|w|^{2}\geq{2}{c^{-2}}\}$. Choose some point $q_{0}\in\Omega^{\prime}$. Since $\Omega\setminus A_{p}$ is path-connected, one can choose a path $\gamma$ that connects $q_{0}$ and $q$. Suppose under $T$ the image of $\gamma$ intersects $\partial\mathbb{B}^{n}$ firstly at some point $T(q_{1})$.
Along the path $\gamma$ take a sequence of points $(q_{l})_{l\in\mathbb{N}}\subset\Omega^{\prime}$ such that $q_{l}\to q_{1}$. Then by (3.5),
$\Phi_{p}(q_{l})={\frac{-2}{c^{2}}}\log\left(1-\frac{c^{2}}{2}|T(q_{l})|^{2}\right).$
Here, as $q_{l}\to q_{1}$, the left hand side is finite but the right hand side blows up to infinity. This is a contradiction, so we have thus proved our claim, which says that $\Omega^{\prime}=\Omega\setminus A_{p}$. Therefore, (3.5) in fact holds on $\Omega\setminus A_{p}$.
2) Since $T$ maps $\Omega\setminus A_{p}$ to the ball $\mathbb{B}^{n}$ and satisfies
$|T(z)|^{2}<{2}{c^{-2}},$ (3.6)
by the Riemann removable singularity theorem, $T$ extends across the analytic variety $A_{p}$ to the whole domain $\Omega$ with $|T(z)|^{2}\leq{2}{c^{-2}}.$ The maximum modulus principle yields that (3.6) in fact holds on $\Omega$.
Lastly, for general $p\in U$, one performs a possible linear transformation $F$ from $\Omega$ to $\Omega_{1}$ such that the Bergman metric on $\Omega_{1}$ satisfies (3.3) at $F(p)$. Since $F$ is a biholomorphism, the Bergman metric on $\Omega_{1}$ locally has the same holomorphic sectional curvature.
Denote by $\omega_{\mathcal{B}}$ the Bergman metric on $\mathcal{B}$. From the above proof, we see that the Kähler potential for the Bergman metric has the following formula.
$\Phi_{p}(z)={\frac{-2}{c^{2}}}\log\left(1-\frac{c^{2}}{2}\sum_{\alpha,\beta=1}^{n}w_{\alpha}(z)g_{\alpha\bar{\beta}}(p)\overline{w_{\beta}(z)}\right),\quad z\in\Omega.$
Direct computations then yield that $T:\Omega\to\mathcal{B}$ is an isometry. In fact, if the Bergman metric at $p\in U$ satisfies (3.3), then the potential $\Phi_{p}$ simplifies to
$\Phi_{p}(z)={\frac{-2}{c^{2}}}\log\left(1-\frac{c^{2}}{2}|T(z)|^{2}\right),\quad z\in\Omega.$
So,
$g_{\alpha\bar{\beta}}(z)$ $=\frac{\partial^{2}\Phi_{p}(z)}{\partial z_{\alpha}\partial\overline{z_{\beta}}}$
$=\sum_{i,j=1}^{n}\left(1-\frac{c^{2}}{2}|T(z)|^{2}\right)^{-2}\left[\delta_{ij}\left(1-\frac{c^{2}}{2}|T(z)|^{2}\right)+\frac{c^{2}}{2}\overline{w_{i}(z)}w_{j}(z)\right]{\frac{\partial w_{i}(z)}{\partial{z_{\alpha}}}}\frac{\partial\overline{w_{j}(z)}}{\partial{\overline{z_{\beta}}}}$
$=\sum_{i,j=1}^{n}G_{i\bar{j}}(w){\frac{\partial w_{i}(z)}{\partial{z_{\alpha}}}}\frac{\partial\overline{w_{j}(z)}}{\partial{\overline{z_{\beta}}}},\quad z\in\Omega,$
where
$G_{i\bar{j}}(w):=\frac{\partial^{2}}{\partial w_{i}\partial\overline{w_{j}}}\left(\frac{-2}{c^{2}}\log(1-\frac{c^{2}}{2}|w|^{2})\right)$
gives the Bergman metric on $\mathbb{B}^{n}$. The general case follows in a similar manner. And the curvature part also follows.
Theorem 3.4 directly implies a stronger version of Lu’s classical theorem (Theorem 1.4), requiring constant curvature only on an open set rather than globally. This insight, combined with our study of the Lu constant, provides the key to completing the proof of Theorem 1.10.
Proof of Theorem 1.10.
From the upper bound of the holomorphic sectional curvature of the Bergman metric and the Ahlfors-Schwarz lemma (see [2, 42, 49]), one obtains the first inequality in the following inequality chain
$K_{\Omega}\geq L(\Omega)B_{\Omega}\geq C_{\Omega},$
where the second inequality holds true due to the definition of $L(\Omega)$. By Theorem 2.2, there exists some $\epsilon>0$ such that: if $z\in\Omega$ and $\delta_{\Omega}(z)<\epsilon$, then there is a nonempty, open set $E_{\Omega}(z)$ of directions where the Kobayashi metric $K_{\Omega}$ and the Carathéodory metric $C_{\Omega}$ coincide. One gets on $E_{\Omega}(z)$ that
$K_{\Omega}=L(\Omega)B_{\Omega}=C_{\Omega}.$
Using the argument of the third author in [47], we know that the holomorphic sectional curvature $H(z;v)$ of the Bergman metric is equal to $-2(L(\Omega))^{2}$ for any non-zero $v\in E_{\Omega}(z)$. Since the holomorphic sectional curvature function
$v\in T_{z}(\Omega)\mapsto H(z;v)\in\mathbb{R}$
is algebraic ($z$ is fixed), we obtain that the holomorphic sectional curvature $H(z;v)$ is equal to $-2(L(\Omega))^{2}$ for all non-zero vector $v\in T_{z}(\Omega)$. This is true for all $z\in\Omega$ such that $\delta_{\Omega}(z)<\epsilon$ for a sufficiently small $\epsilon$. Then, the holomorphic sectional curvature of the Bergman metric $B^{2}_{\Omega}$ is equal to $-2(L(\Omega))^{2}$ on $\Omega\setminus\Sigma$, where $\Sigma$ is defined by (2.1). Since the Bergman metric $(\Omega,B^{2}_{\Omega})$ is complete, Theorem 1.4 yields that $\Omega$ must be biholomorphic to $\mathbb{B}^{n}$.
Statements and Declarations
No financial or non-financial interests that are directly or indirectly related to the work submitted for publication were reported by the authors.
Data sharing is not applicable to this article as no datasets were generated or analysed during the current study.
Funding
The research of the first author was partially supported by an AMS-Simons travel grant and the NSF grant DMS-2103608.
Acknowledgements
The authors sincerely thank Łukasz Kosiński, Bernhard Lamel, Damin Wu and Andrew Zimmer for their comments. The first author thanks Song-Ying Li, Ming Xiao, and Hang Xu for stimulating discussions.